A firm that has a total cost function TC(Q) 7Q 2 exhibits: Wählen Sie eine Antwort: O a. Constant returns of scale b. Decreasing returns of scale O C. Economies of scale O d. Diseconomies of scale
Answer #1

TC = 7Q - 2

ATC = TC/Q = 7 - 2/Q

ATC = 7 - 2/Q

As we increase Q, the ATC will increase

So this firm will have diseconomies of scale.

As Q rises, the term 2/Q will fall and it is a negative term, total ATC will rise as Q rises,

So correct ans is d) diseconomies of scale.